OpenStudy (toxicsugar22):

hi can u help me solve this 3^(9-3x)=27

OpenStudy (watchmath):

nobody help you?

OpenStudy (toxicsugar22):

no can u plese

OpenStudy (watchmath):

write 27 using base 3 as $3^3$ Compare the exponents 9-3x=3 -3x=-6 x=2

OpenStudy (toxicsugar22):

wait could u tell me what happened to the 27 pleae

OpenStudy (toxicsugar22):

in this equation im still confused i have a test tommorow coul u give me some promblem on this so i know how to do it

OpenStudy (toxicsugar22):

hello are u there

OpenStudy (angela210793):

3^(9-3x)=27 3^(9-3x)=3^3 since the base is the same=3 we can equal the exponents as well --> 9-3x=3 -3x=3=9 -3x=-6 3x=6 x=6/3 x=2